How many ounces of pure copper must a goldsmith add to 240 ounces of an alloy containing 40% copper to obtain an alloy containing 60% copper?

Answer by checkley79(3341)

.40*240+X=.60(240+X)
96+X=144+.60X
X-.60X=144-96
.40X=48
X=48/.40
X=120 OZ OF PURE COPPER IS NEEDED.
PROOF:
.40*240+120=.60(240+120)
96+120=.60*360
216=216